I always have problems with word problems, Can you pleae help with this one

A rectangular painting with a width of
x centimeters has an area of x2+50x square centimeters.
Find a binomial that represents the length

Answer by RAY100(1637)

Area = Length * Width
.
(x^2 +50x) = L*x
.
x(x+50) = L*x
.
(x+50) = L
.
check...A=LW,,,A=(x+50)(x) = x^2 +50x,,,,ok